A certain electromagnetic wave has a wavelength of 425 nm. What is the frequency of the wave? What region of the electromagnetic spectrum is this? What is the energy of the wave in Joules?
1
Expert's answer
2019-10-01T03:24:12-0400

Frequency: (3*10^8)/(4.25*10^(-7))=7.06*10^14

UV-part of the spectrum

Ebergy: 6.63*10^(-34)*7.06*10^14=4.68*10^(-19)